# arXiv:2603.01466v1[math.FA] 02 Mar 2026

### v1

@@ 4. Maximal violation of bilocal inequalities and algebraic structures
Proof. We first deal with the special case: $\mathcal{M}_{A}=\mathcal{R}_{A}$ and $\mathcal{M}_{C}=\mathcal{R}_{C}$. Let $\widetilde{\mathcal{H}}=\otimes_{\alpha\in\mathbb{N}}(\mathcal{H}_{\alpha},\Omega_{\alpha})$ be the incomplete tensor product of $\mathcal{H}_{\alpha}=\mathbb{C}^{2}\otimes\mathbb{C}^{2},\ (\alpha\in\mathbb{N})$, with $\Omega_{\alpha}=\frac{1}{\sqrt{2}}(\Phi_{1}\otimes\Phi_{1}+\Phi_{2}\otimes\Phi_{2})\equiv\Omega\text{ for some basis }\{\Phi_{1},\Phi_{2}\}\in\mathbb{C}^{2},\alpha\in\mathbb{N}.$ Then by the construction of Araki and Woods [76], $\mathcal{R}_{A}\simeq\mathcal{R}(\mathcal{H}_{\alpha},\mathcal{B}(\mathbb{C}^{2})\otimes 1,\Omega_{\alpha},\alpha\in\mathbb{N})$ holds, using their notation for the factor on the right-hand side which is the infinite tensor product of $\mathcal{B}(\mathbb{C}^{2})\otimes 1$ with itself on $\widetilde{\mathcal{H}}$. Similarly, let $\widetilde{\mathcal{K}}=\otimes_{\alpha\in\mathbb{N}}(\mathcal{K}_{\alpha},\Lambda_{\alpha})$ be the incomplete tensor product of $\mathcal{K}_{\alpha}=\mathbb{C}^{2}\otimes\mathbb{C}^{2},\ (\alpha\in\mathbb{N})$, with $\Lambda_{\alpha}=\frac{1}{\sqrt{2}}(\Psi_{3}\otimes\Psi_{3}+\Psi_{4}\otimes\Psi_{4})\equiv\Lambda\text{ for some basis }\{\Psi_{3},\Psi_{4}\}\in\mathbb{C}^{2},\alpha\in\mathbb{N}.$ Then $\mathcal{R}_{C}\simeq\mathcal{R}(\mathcal{K}_{\alpha},\mathcal{B}(\mathbb{C}^{2})\otimes 1,\Lambda_{\alpha},\alpha\in\mathbb{N})$ holds, using their notation for the factor on the right-hand side which is the infinite tensor product of $\mathcal{B}(\mathbb{C}^{2})\otimes 1$ with itself on $\widetilde{\mathcal{K}}$. We can obtain that there is $\mathcal{R}_{C}\simeq\mathcal{R}(\mathcal{K}_{\alpha},\mathcal{B}(\mathbb{C}^{2})\otimes 1,\Lambda_{\alpha},\alpha\in\mathbb{N})$. Now from Theorem 5, we know that $I,\ a_{0},\ a_{1}$ and $a_{2}=-\frac{i}{2}[a_{0},a_{1}]$ form a realization of the Pauli spin matrices on $\mathcal{H}_{\alpha}$, and the same holds for $I,\ c_{0},\ c_{1}$ and $c_{2}=-\frac{i}{2}[c_{0},c_{1}]$ in $\mathcal{K}_{\alpha}$, and $b_{0},b_{1}$ with $b_{i}^{2}=I$ such that the quantity (3.1) arrives at $2\sqrt{2}$. Applying the similar technologies of Eqs. (6)-(8) of Ref.[72], one can obtain that
@@
(4.10) $\sqrt{|\langle\Omega\otimes\Lambda,[(a_{0}-a_{1})b_{1}(c_{0}-c_{1})]\Omega\otimes\Lambda\rangle|}=\sqrt{2}$
@@
and furthermore $\Phi=\bigotimes_{\alpha\in\mathbb{N}}\Phi_{\alpha},\Psi=\bigotimes_{\alpha\in\mathbb{N}}\Psi_{\alpha},$ with $\Phi_{\alpha}=\Omega_{\alpha}\otimes\Omega_{\alpha}$ and $\Psi_{\alpha}=\Lambda_{\alpha}\otimes\Lambda_{\alpha}$ for almost all $\alpha$. By definition of the incomplete infinite tensor product, such vectors are total in $\widetilde{\mathcal{H}}\otimes\widetilde{\mathcal{K}}$. But one has
$\sqrt{\langle\Phi,\widetilde{T}_{\alpha}\Psi\rangle}=\langle\Phi,\Psi\rangle\cdot\sqrt{2},\sqrt{\langle\Phi,\widetilde{S}_{\alpha}\Psi\rangle}=\langle\Phi,\Psi\rangle\cdot\sqrt{2}.$
for all $\alpha$ larger than some $\alpha_{0}$ (depending on $\Phi$ and $\Psi$).
@@
It is mentioned that the results in Theorem 6 can be applied into the quantum field theory. In the algebraic framework of the QFT, wedge algebras (von Neumann algebras associated to specific unbounded regions in Minkowski spacetime, such as $\mathcal{O}$ associated to tangent regions are typically type III1 factors (a classification indicating the algebra has no trace and no pure normal states, reflecting the infinite degrees of freedom of QFT). By the classical result of A. Connes, all injective (i.e., amenable or hyperfinite, a property related to physical phase space compactness) infinite factors (excepting some type III0) are strongly stable (meaning they can absorb a hyperfinite II1 factor: $\mathcal{M}\simeq\mathcal{M}\otimes\mathcal{R}_{1}$). So the assumption of Theorem 6 is satisfied.
Corollary 7. Let $\mathcal{O}_{i}$ $(i=1,2,3)$ be wedge-shaped regions, $\mathcal{A}(\mathcal{O}_{i})$ the wedge algebra on a separable Hilbert space $H$ with cyclic and separating vector. If each $\mathcal{A}(\mathcal{O}_{i})$ is III1 factor, then $\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=2\sqrt{2}$ for every normal state $\tau$ on $(\mathcal{A}(\mathcal{O}_{1})\vee\mathcal{A}(\mathcal{O}_{2})\vee\mathcal{A}(\mathcal{O}_{3}))^{\prime\prime}$.
Acknowledgement. All authors contributed equally to this work.

### v2

@@ 5. Maximal violation of bilocal inequalities and algebraic structures
Proof. We first deal with the special case: $\mathcal{M}_{\mathcal{A}}=\mathcal{R}_{\mathcal{A}}$ and $\mathcal{M}_{\mathcal{C}}=\mathcal{R}_{\mathcal{C}}$. Let $\widetilde{\mathcal{H}}=\otimes_{\alpha\in\mathbb{N}}(\mathcal{H}_{\alpha},\Omega_{\alpha})$ be the incomplete tensor product of $\mathcal{H}_{\alpha}=\mathbb{C}^{2}\otimes\mathbb{C}^{2},\ (\alpha\in\mathbb{N})$, with $\Omega_{\alpha}=\frac{1}{\sqrt{2}}(\Phi_{1}\otimes\Phi_{1}+\Phi_{2}\otimes\Phi_{2})\equiv\Omega\text{ for some basis }\{\Phi_{1},\Phi_{2}\}\in\mathbb{C}^{2},\alpha\in\mathbb{N}.$ Then by the construction of Araki and Woods [79], $\mathcal{R}_{\mathcal{A}}\simeq\mathcal{R}(\mathcal{H}_{\alpha},\mathcal{B}(\mathbb{C}^{2})\otimes 1,\Omega_{\alpha},\alpha\in\mathbb{N})$ holds, using their notation for the factor on the right-hand side which is the infinite tensor product of $\mathcal{B}(\mathbb{C}^{2})\otimes 1$ with itself on $\widetilde{\mathcal{H}}$. Similarly, let $\widetilde{\mathcal{K}}=\otimes_{\alpha\in\mathbb{N}}(\mathcal{K}_{\alpha},\Lambda_{\alpha})$ be the incomplete tensor product of $\mathcal{K}_{\alpha}=\mathbb{C}^{2}\otimes\mathbb{C}^{2},\ (\alpha\in\mathbb{N})$, with $\Lambda_{\alpha}=\frac{1}{\sqrt{2}}(\Psi_{3}\otimes\Psi_{3}+\Psi_{4}\otimes\Psi_{4})\equiv\Lambda\text{ for some basis }\{\Psi_{3},\Psi_{4}\}\in\mathbb{C}^{2},\alpha\in\mathbb{N}.$ Then $\mathcal{R}_{\mathcal{C}}\simeq\mathcal{R}(\mathcal{K}_{\alpha},\mathcal{B}(\mathbb{C}^{2})\otimes 1,\Lambda_{\alpha},\alpha\in\mathbb{N})$ holds, using their notation for the factor on the right-hand side which is the infinite tensor product of $\mathcal{B}(\mathbb{C}^{2})\otimes 1$ with itself on $\widetilde{\mathcal{K}}$. We can obtain that there is $\mathcal{R}_{\mathcal{C}}\simeq\mathcal{R}(\mathcal{K}_{\alpha},\mathcal{B}(\mathbb{C}^{2})\otimes 1,\Lambda_{\alpha},\alpha\in\mathbb{N})$. Now, from Theorem 5, we know that $A_{0},\ A_{1}$ and $A_{2}=-\frac{i}{2}[A_{0},A_{1}]$ form a realization of the Pauli spin matrices on $\mathcal{H}_{\alpha}$, and the same holds for $C_{0},\ C_{1}$ and $C_{2}=-\frac{i}{2}[C_{0},C_{1}]$ in $\mathcal{K}_{\alpha}$, and $B_{0},B_{1}$ with $B_{i}^{2}=I$ such that the quantity $\mathcal{S}_{\tau}$ in Ineq. (4.1) arrives at $2\sqrt{2}$. Applying similar technologies of Eqs. (6)-(8) of Ref.[77], one can obtain that
@@
(5.10) $\sqrt{|\langle\Omega\otimes\Lambda,[(A_{0}-A_{1})B_{1}(C_{0}-C_{1})]\Omega\otimes\Lambda\rangle|}=\sqrt{2}.$
@@
Furthermore, $\Xi=\bigotimes_{\alpha\in\mathbb{N}}\Xi_{\alpha},\Upsilon=\bigotimes_{\alpha\in\mathbb{N}}\Upsilon_{\alpha},$ with $\Xi_{\alpha}=\Omega_{\alpha}\otimes\Omega_{\alpha}$ and $\Upsilon_{\alpha}=\Lambda_{\alpha}\otimes\Lambda_{\alpha}$ for almost all $\alpha$. By the definition of the incomplete infinite tensor product, such vectors are total in $\widetilde{\mathcal{H}}\otimes\widetilde{\mathcal{K}}$. But one has
$\sqrt{\langle\Xi,\widetilde{T}_{\alpha}\Upsilon\rangle}=\langle\Xi,\Upsilon\rangle\cdot\sqrt{2},\sqrt{\langle\Xi,\widetilde{S}_{\alpha}\Upsilon\rangle}=\langle\Xi,\Upsilon\rangle\cdot\sqrt{2}.$
for all $\alpha$ larger than some $\alpha_{0}$ (depending on $\Xi$ and $\Upsilon$).
@@
It is mentioned that the results of Theorem 6 can be applied to quantum field theory. In the algebraic framework of the QFT, wedge algebras (von Neumann algebras associated to specific unbounded regions in Minkowski spacetime, such as $\mathcal{O}$ associated to tangent regions, are typically type III1 factors. By the classical result of A. Connes, all injective infinite factors (except some type III0) are strongly stable (meaning they can absorb a hyperfinite II1 factor: $\mathcal{M}\simeq\mathcal{M}\overline{\otimes}\mathcal{R}_{1}$).
Corollary 7. Let $\mathcal{O}_{i}$ $(i=1,2,3)$ be wedge-shaped regions, $\mathcal{A}(\mathcal{O}_{i})$ be the wedge algebra on a separable Hilbert space $H$ with cyclic and separating vector. If $\mathcal{A}(\mathcal{O}_{1})$ and $\mathcal{A}(\mathcal{O}_{3})$ are both type III1 factors, then $\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=2\sqrt{2}$ for every normal state $\tau$ on $(\mathcal{A}(\mathcal{O}_{1})\vee\mathcal{A}(\mathcal{O}_{2})\vee\mathcal{A}(\mathcal{O}_{3}))^{\prime\prime}$.
Acknowledgement. Authors thank for your comments. This work is supported by the National Natural Science Foundation of China (Grant No. 12271394).